# Existence for Stable Rotating Star-Planet Systems

### v1

@@ 7 Upper Bounds for Distances Between Connected Components of Minimizers
We put the condition $\gamma>3/2$ so that we can apply Theorem 2.13, but in fact the estimate (48) actually holds even for smaller values of $\gamma$.
Actually, similar result can be shown for stars’ densities.
@@
Given polytropic law $P(\rho)=K\rho^{\gamma}$ indexed by $\gamma>\frac{3}{2}$ and a constrained minimizer $\rho(m)=\rho_{1-m}+\rho_{m}$ mentioned in Theorem 2.13. There is a $C>0$ and a $\delta>0$, such that for all $m\in(0,\delta)$, if spt $\rho_{1-m}$ has 2 convex connected components, then the distance between these 2 components is less then $\eta$, where $\eta=Cm^{5}$.
@@
$\begin{split}&\quad E_{J}\left({\rho(m)}\right)-E_{J}\left({\rho^{[h]}(m)}\right)\\ &\geq\left({\frac{1}{R^{3}\left({1-m}\right)}-\frac{J^{2}}{I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\frac{1}{1-m}}\right){\int_{\{{\rho_{U}{(x)}>0,\rho_{V}{(y)}>0}\}}{\rho_{U}(x)\rho_{V}(y)\left({y_{1}-x_{1}}\right)hd^{3}xd^{3}y}}\\ &\quad+O\left(h^{2}\right)-\frac{16m_{1}m_{2}m^{5}\left({1-m}\right)^{3}}{J^{4}}h\\ &\geq\left({\frac{dm_{1}m_{2}}{2R^{3}\left({1-m}\right)}-\frac{16m_{1}m_{2}m^{5}\left({1-m}\right)^{3}}{J^{4}}}\right)h+O\left(h^{2}\right)\end{split}$
We just note that for star density, R(1m)R(1-m) will bounded by $R(J)$ as $m$ goes to 0 ($R(J)$ is given in Corollary 5.26). Thus we know $\frac{1}{R^{3}\left({1-m}\right)}\geq\frac{2J^{2}}{I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\frac{1}{1-m}$ for $m$ small enough. Moreover, when $d>\frac{32m^{5}\left({1-m}\right)^{3}R^{3}(1-m)}{J^{4}}$, we know $E_{J}\left({\rho(m)}\right)-E_{J}\left({\rho^{[h]}(m)}\right)>0$ for $h$ small enough. Thus we can apply the method of contradiction to prove this corollary. ∎
@@
We have estimated distance between (possible) two connected components. For cases with more than two connected components, say $U_{1},U_{1},U_{3},$... (could be uncountable many), if they satisfy certain shape configurations — for example $U_{i}$ and $V_{i}$ satisfy (42) up to rotation, where $V_{i}={\bigcup\limits_{\alpha\neq i}U_{\alpha}}$ — we can then move $U_{i}$ and $V_{i}$ similarly as above and analyze the distance relationships between them. Furthermore, we can also use Theorem 2.13 (ii) to rule out some configurations of connected components. For example, the case where an outer shell encloses an inner solid sphere is impossible for $\rho_{m}$ or $\rho_{1-m}$.
@@
It should be noted that the troublesome term in (48) preventing us from further reducing the distance between connected components $U$ and $V$ is the term $\frac{16m_{1}m_{2}m^{3}\left({1-m}\right)^{5}}{J^{4}}$, which comes from (47). In other words, we might be able to prove simple connectedness in each subdomain by moving and gluing the relevant connected components, if we can show that $G\left({\rho_{m}^{[h]},\rho_{m}^{[h]}}\right)-G\left({\rho_{m},\rho_{m}}\right)$ can absorb $\left|{G\left({\rho_{1-m},\rho_{m}-\rho_{m}^{[h]}}\right)}\right|$. For instance, this could be achieved by proving $G\left({\rho_{m}^{[h]},\rho_{m}^{[h]}}\right)-G\left({\rho_{m},\rho_{m}}\right)>10\left|{G\left({\rho_{1-m},\rho_{m}-\rho_{m}^{[h]}}\right)}\right|$ for all small $m$ (that is $m\in(0,\delta)$ for some $\delta>0$) and for all positive distance $d$ between two connected components $U$ and $V$.
@@
The support of $\rho(m)$ in Theorem 2.13 should have exactly 2 connected components when $m$ small enough.
@@
Appendix A Appendix
Here we introduce some preliminary knowledge which are used in the paper.
A.1 Calculus of Variations in Vector Spaces
Since the functional we deal with is not differentiable in any usual sense, we begin with a less restrictive notion of differentiability.
Let $X$ be a real vector space, $U$ a subset of $X$, and $E$ an extended real-valued function defined on $U$. Given $u\in U$, let $P(u)=\{v\in X\mid\exists\epsilon(v)>0,\text{ such that }u+tv\in U\text{ for }t\in[0,\epsilon(v))\}$. This set is a cone in $X$. If $P$ is any cone in $X$, let $X(P)$ be the linear subspace of $X$ generated by $P$. We say that $E$ is $P$-differentiable at $u$ if $P\subset P(u)$ and there is a linear functional $\mu$ on $X(P)$ such that $\forall v\in P$
$\lim\limits_{t\rightarrow 0^{+}}t^{-1}\{E(u+tv)-E(u)\}=\mu(v),$
The linearity of $\mu$ is needed in our analysis; its precise domain is not essential here. We denote the “derivative” $\mu$ by $E_{u}^{\prime}$—with $P$ understood from context.
We say $E$ has a weak local minimum at $u\in U$ if for each $v\in P(u)$, there exists $\{t_{n}\}$ such that $t_{n}>0$, ${\lim\limits_{n\rightarrow\infty}t_{n}}=0$ and $E\left({u+t_{n}v}\right)\geq E(u)$ for $n$ large enough. And $u$ is called weak local minimizer. We discuss more about the motivation of weak local minimizer in [15].
Let $W\subset\left\{u\in U\middle|g(u)=M\right\}$, where $g$ is a linear functional on $X$ and $M$ is a constant. Then similarly, we say $E$ has a weak local minimum at $u$ restricted to $W$ if $u\in W$ and for each $v\in P(u)$, there exists $t_{n}$} such that $t_{n}>0$, ${\lim\limits_{n\rightarrow\infty}t_{n}}=0$, and for for $n$ large enough, we have $u+t_{n}v\in W$, and $E\left({u+t_{n}w}\right)\geq E(u)$. Furthermore, we have a result for weak local minimum restricted to $W$:
Theorem A.1 ((Generalized) Lagrange Multiplier Theorem [4, Section 2] [15, Appendix A]).
Suppose $E$ is $P$-differentiable at $u\in U$ and that its restriction to $W$ has a weak local minimum at $u$. Suppose also that $P$ is convex and that there is a $u_{0}\in P$ such that $u_{0}\in P$ and $g\left(u_{0}\right)\neq 0$. Then there is a unique constant $\lambda$ such that $E_{u}^{\prime}(v)\geq\lambda g(v)$, all $v\in P$.
A.2 Properties of Sobolev Spaces
We recall some properties of Sobolev spaces as well as of $L^{p}$ spaces. Some of them will be used frequently in this paper.
Proposition A.2 (Interpolation Inequality [8, Section 4.2]).
If $f\in L^{p}\cap L^{q}$ with $1\leq p\leq q\leq\infty$, then $f\in L^{r}$ for all $r,p\leq r\leq q$, and the following interpolation inequality holds:
$\|f\|_{r}\leq\|f\|_{p}^{\alpha}\|f\|_{q}^{1-\alpha},\text{ where }\frac{1}{r}=\frac{\alpha}{p}+\frac{1-\alpha}{q},0\leq\alpha\leq 1$
@@ A.2 Properties of Sobolev Spaces
It turns out for potential $V_{\rho}$ given in (9), we have $\frac{\partial V_{\rho}}{\partial x_{j}}(x)=-\int\frac{y_{j}\rho(x-y)}{|y|^{3}}d^{3}y$ not only in the distribution sense but also in the classical sense:
Proposition A.4 (Differentiability of Potential [4, Proposition 7] [15, Proposition A.13]).
@@
To prove the low semicontinuity of the energy, we hope to connect the weak convergence of densities and strong convergence of potentials.
Proposition A.5 (Compactness of Convolution (Potential) Operator with $q\geq 1$ [16, Proposition A.7]).
Let $G(x)=\frac{1}{|x|}$, then $G\in L_{w}^{3}\left(\mathbb{R}^{3}\right)$. Let F=GBF=G*B, where $B$ is a bounded set in $L^{q}\left(\mathbb{R}^{3}\right)$ with $q\geq 1$. Then let $\Omega\subset\mathbb{R}^{3}$ is open, bounded and of class $C^{1}$.
  1. (1)
    If $1\leq q<\frac{3}{2}$, then $\forall 1\leq r<p$, where $\frac{1}{p}=\frac{1}{q}-\frac{2}{3},\left.F\right|_{\Omega}$ has compact closure in $L^{r}(\Omega)$.
  2. (2)
    If $q=\frac{3}{2}$, we assume in addition $B$ is bounded in $L^{1}\left(\mathbb{R}^{3}\right)$, then $\forall 1\leq r<\infty,\left.F\right|_{\Omega}$ has compact closure in $L^{r}(\Omega)$.
  3. (3)
    If $q>\frac{3}{2}$, we assume in addition $B$ is bounded in $L^{1}\left(\mathbb{R}^{3}\right)$, then $\left.F\right|_{\Omega}$ has compact closure in $C(\bar{\Omega})$.
@@ Acknowledgments
The author is partially supported by the National Science Foundation grant DMS-2308208. This work was primarily carried out during the author’s Master’s studies at the University of Bonn. The author thanks Juan Velázquez (master’s thesis advisor) and Dimitri Cobb (second advisor). Thanks also to Lorenzo Pompili, Shao Liu, Xiaopeng Cheng, Bernhard Kepka, Daniel Sánchez Simón del Pino, Han Cao for discussions, and to Théophile Dolmaire and other instructors. The author is grateful to his parents.

### v2

@@ 7 Upper Bounds for Distances Between Connected Components of Minimizers
We put the condition $\gamma>\frac{3}{2}$ so that we can apply Theorem 2.13, but in fact the estimate (49) actually holds even for smaller values of $\gamma$.
Actually, a similar result can be shown for stars’ densities.
@@
Given polytropic law $P(\rho)=K\rho^{\gamma}$ indexed by $\gamma>\frac{3}{2}$ and a constrained minimizer $\rho(m)=\rho_{1-m}+\rho_{m}$ mentioned in Theorem 2.13. There is a $C>0$ and a $\delta>0$, such that for all $m\in(0,\delta)$, if spt $\rho_{1-m}$ has 2 convex connected components, then the distance between these 2 components is less than $\eta$, where $\eta=Cm^{5}$.
@@
$\begin{split}&\quad E_{J}\left({\rho(m)}\right)-E_{J}\left({\rho^{[h]}(m)}\right)\\ &\geq\left({\frac{1}{R^{3}\left({1-m}\right)}-\frac{J^{2}}{I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\frac{1}{1-m}}\right){\iint_{\{{\rho_{U}{(x)}>0,\rho_{V}{(y)}>0}\}}{\rho_{U}(x)\rho_{V}(y)\left({y_{1}-x_{1}}\right)h\,dx\,dy}}\\ &\quad+O\left(h^{2}\right)-\frac{16m_{1}m_{2}m^{5}\left({1-m}\right)^{3}}{J^{4}}h\\ &\geq\left({\frac{dm_{1}m_{2}}{2R^{3}\left({1-m}\right)}-\frac{16m_{1}m_{2}m^{5}\left({1-m}\right)^{3}}{J^{4}}}\right)h+O\left(h^{2}\right)\end{split}$
We just note that for star density, R(1m)R(1-m) will be bounded by $R(J)$ as $m$ goes to 0 ($R(J)$ is given in Corollary 5.25). Thus we know $\frac{1}{R^{3}\left({1-m}\right)}\geq\frac{2J^{2}}{I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\frac{1}{1-m}$ for $m$ small enough. Moreover, when $d>\frac{32m^{5}\left({1-m}\right)^{3}R^{3}(1-m)}{J^{4}}$, we know $E_{J}\left({\rho(m)}\right)-E_{J}\left({\rho^{[h]}(m)}\right)>0$ for $h$ small enough. Thus we can apply the method of contradiction to prove this corollary. ∎
@@
We have estimated distance between (possible) two connected components. For cases with more than two connected components, say $U_{1},U_{1},U_{3},$... (could be uncountable many), if they satisfy certain shape configurations — for example $U_{i}$ and $V_{i}$ satisfy (43) up to rotation, where $V_{i}={\bigcup\limits_{\alpha\neq i}U_{\alpha}}$ — we can then move $U_{i}$ and $V_{i}$ similarly as above and analyze the distance relationships between them. Furthermore, we can also use Theorem 2.13 (ii) to rule out some configurations of connected components. For example, the case in which an outer shell encloses an inner solid sphere is impossible for $\rho_{m}$ or $\rho_{1-m}$.
@@
It should be noted that the troublesome term in (49) preventing us from further reducing the distance between connected components $U$ and $V$ is the term $\frac{16m_{1}m_{2}m^{3}\left({1-m}\right)^{5}}{J^{4}}$, which comes from (48). In other words, it may become possible to prove simple connectedness in each subdomain by moving and gluing the relevant connected components, if one can show that $G\left({\rho_{m}^{[h]},\rho_{m}^{[h]}}\right)-G\left({\rho_{m},\rho_{m}}\right)$ can absorb $\left|{G\left({\rho_{1-m},\rho_{m}-\rho_{m}^{[h]}}\right)}\right|$. For instance, this could be achieved by proving $G\left({\rho_{m}^{[h]},\rho_{m}^{[h]}}\right)-G\left({\rho_{m},\rho_{m}}\right)>10\left|{G\left({\rho_{1-m},\rho_{m}-\rho_{m}^{[h]}}\right)}\right|$ for all small $m$ (that is $m\in(0,\delta)$ for some $\delta>0$) and for all positive distance $d$ between two connected components $U$ and $V$.
@@
The support of $\rho(m)$ in Theorem 2.13 should have exactly 2 connected components when $m$ is small enough.
@@
Appendix
Appendix A Properties of Sobolev Spaces
Here we recall two propositions about the potential $V_{\rho}(x)=\int_{\mathbb{R}^{3}}\frac{\rho(y)}{|x-y|}\,dy=\left(\frac{1}{|\cdot|}*\rho\right)(x)$.
@@ Appendix A Properties of Sobolev Spaces
Proposition A.2 (Differentiability of Potential [4, Section 3] [10, Appendix B]).
@@ Acknowledgments
The author is partially supported by the National Science Foundation grant DMS-2308208. This work was primarily carried out during the author’s Master’s studies at the University of Bonn. The author thanks Juan Velázquez and Dimitri Cobb for their continued advice and support since the author’s time in Bonn, as well as Christof Sparber and Mimi Dai for their comments and support during the author’s Ph.D. studies at the University of Illinois Chicago. Thanks also to Lorenzo Pompili, Shao Liu, Xiaopeng Cheng, Bernhard Kepka, Daniel Sánchez Simón del Pino, Han Cao for discussions, and to Théophile Dolmaire and other instructors. The author is grateful to his parents.